A car with a mass of 2500 kg and moving at 3 m/s northeast collides with a 3000 kg car moving northwest at 2.5 m/s . Find their final velocities. Assume a perfectly inelastic collision.



please answer!!!



1
Expert's answer
2021-03-23T11:13:08-0400

"p_{1x}+p_{2x}=p_x\to p_x=2500\cdot3\cdot\cos45°-3000\cdot2.5\cdot\cos45°=0"


"p_{1y}+p_{2y}=p_y\to p_y=2500\cdot3\cdot\sin45°+3000\cdot2.5\cdot\sin45°=10607\ (kg\cdot m/s)"


"p=\sqrt{p_x^2+p_y^2}=\sqrt{0^2+10607^2}=10607\ (kg\cdot m/s)"


"p=(m_1+m_2)v\to v=p/(m_1+m_2)=10607/(2500+3000)=1.93\ (m/s)" . Answer